16. Find the value of n such that x2 + 2x + n is a perfect square trinomial.
4
–1
1
2

Answer by Theo(13342)

based on the attached perfect square trinomial tutorial, I believe your answer would be:

x^2 + 2x + 1 is a perfect square trinomial.

this can be factored as (x+1)^2 = x^2 + 2x + 1

you take 1/2 the b term and square it to get the c term.

in your equation, 2 is the b term. 1/2 of 2 is 1. 1^2 is 1.

standard form of this equation is:

ax^2 + bx + c where a = coefficient of the x^2 term, b = coefficient of the x term, c = constant.